OpenStudy (lakers):

If the area of a rhombus is 126 m2, and one diagonal measures 18 m, find the length of the other diagonal. A. 14 cm B. 7 cm C. 3.5 cm D. 72 cm

OpenStudy (anonymous):

A.14 cm

OpenStudy (lakers):

why not B?

OpenStudy (anonymous):

why is it 14 cm?

OpenStudy (lakers):

@RyanL. its B right?

OpenStudy (anonymous):

Because the S=(d1*d2)/2 so we have 126=(18*X)/2 or 126*2=18*X X=252/18=14

OpenStudy (lakers):

ok i see wat yu mean thanks @olya_ua

OpenStudy (anonymous):

u'r welcome lakers)